The base of a mountain is at sea level where the gravitational field strength is 9.810N/kg. The value of the gravitational field strength at the top of the mountain is 9.790N/kg. Calculate the height of the mountain above the sea level.

1
Expert's answer
2021-02-22T10:27:34-0500

By the definition of the gravitational field strength, we have:


"g=\dfrac{GM}{R^2}."

Then, we can write:


"\dfrac{g_{sea}}{g_{top}}=\dfrac{\dfrac{GM}{R^2}}{\dfrac{GM}{(R+h)^2}}=\dfrac{(R+h)^2}{R^2},""\sqrt{\dfrac{g_{sea}}{g_{top}}}=\dfrac{R+h}{R},""h=R(\sqrt{\dfrac{g_{sea}}{g_{top}}}-1),""h=6.38\cdot10^6\ m\cdot(\sqrt{\dfrac{9.81\ \dfrac{N}{kg}}{9.79\ \dfrac{N}{kg}}}-1)=6513\ m."